# LiDAR Point Cloud-based Multiple Vehicle Tracking with Probabilistic Measurement-Region Association

### v2

@@ Abstract
Multiple extended target tracking (ETT) has gained increasing attention due to the development of high-precision LiDAR and radar sensors in automotive applications. For LiDAR point cloud-based vehicle tracking, this paper presents a probabilistic measurement-region association (PMRA) ETT model, which can describe the complex measurement distribution by partitioning the target extent into different regions. The PMRA model overcomes the drawbacks of previous data-region association (DRA) models by eliminating the approximation error of constrained estimation and using continuous integrals to more reliably calculate the association probabilities. Furthermore, the PMRA model is integrated with the Poisson multi-Bernoulli mixture (PMBM) filter for tracking multiple vehicles. Simulation results illustrate the superior estimation accuracy of the proposed PMRA-PMBM filter in terms of both positions and extents of the vehicles comparing with PMBM filters using the gamma Gaussian inverse Wishart and DRA implementations.
@@ I Introduction
LiDAR and radar point clouds can provide abundant and accurate spatial information of the surrounding environment, which is vital for perception tasks such as target detection and tracking in autonomous driving and intelligent transportation systems [1, 2, 3, 4, 5]. In the context of point cloud-based multiple target tracking (MTT), extended target tracking (ETT) methods have attracted increasing attentions [6, 7, 8]. The ETT differs from traditional MTT approaches by assuming the sensor can gather multiple measurements of a target in each scan, thus allowing for the simultaneous estimation of the target’s location and extent directly from the point cloud [9, 10].
@@
The objective of this study is to overcome the deficiencies of DRA methods and extend the region-partitioning idea to LiDAR point cloud-based multiple vehicle tracking. To this end, we design the probabilistic measurement-region association (PMRA) model, which can directly obtain rectangular extents using the Wishart distribution. The PMRA model also utilizes continuous integrals and the visible angle of regions to improve the accuracy and stability for calculating the association probabilities. For tracking multiple vehicles, we integrate the PMRA model with the Poisson multi-Bernoulli mixture (PMBM) filter, a state-of-the-art MTT framework [15, 16]. Simulation results show that the PMRA-PMBM filter with its particle implementation achieves higher estimation accuracy in terms of both positions and extents of the vehicles compared to GGIW and DRA implementations of the PMBM filter.
@@ II System Modeling
A system model based on random finite set (RFS) and PMBM conjugate prior is applied in this study. At time step $k$, the multi-target state is represented by an RFS $\mathbf{X}_{k}=\{\mathbf{x}_{k}^{i}\}_{i\in\mathbf{I}_{k}}$ where $\mathbf{I}_{k}$ is the target index set with cardinality $|\mathbf{I}_{k}|=N_{k}$; both the single target state $\mathbf{x}_{k}^{i}$ and number of targets $N_{k}$ are random. Similarly, the set of measurements is expressed as $\mathbf{Z}_{k}=\{\mathbf{z}_{k}^{m}\}_{m\in\mathbf{M}_{k}}$ with $|\mathbf{M}_{k}|=M_{k}$, and the collection of all measurement sets from time $1$ to time $k$ is denoted by $\mathbf{Z}^{k}$. For LiDAR-based ETT, the 3D point cloud is usually projected to the 2D bird’s-eye-view (BEV) plane before tracking [6, 7]. In this work, we follow the same approach and define the system model on the global Cartesian coordinate system, as illustrated in Fig. 1.
Figure 1: Illustration of the system model. $(x_{\mathrm{s}},y_{\mathrm{s}})$ and $(x_{\mathrm{t}},y_{\mathrm{t}})$ denote the positions of LiDAR and target. The target extent is divided into five regions $r_{1}$-$r_{5}$. The vertices $\mathbf{p}^{1}$-$\mathbf{p}^{4}$ are determined by the eigenvectors $\{\mathbf{e}^{1},\mathbf{e}^{2}\}$ and eigenvalues $\{e^{1},e^{2}\}$ of the extent matrix $E$.
@@ II-A RFSs and PMBM Posterior Density
The RFSs used in our system modeling are briefly introduced in this section. A Poisson point process (PPP) is an RFS with Poisson distributed cardinality and i.i.d. elements. Thus, the density of PPP RFS is defined by
@@
where $\mu$ is the Poisson rate, $f(\mathbf{x})$ is the spatial density, $D(\mathbf{x})=\mu f(\mathbf{x})$ is the intensity function. A Bernoulli RFS contains a single element with probability $r$ or, is an empty set with probability $1-r$. The density of Bernoulli RFS is
@@
A multi-Bernoulli (MB) RFS is the union of a fixed number of independent Bernoulli RFSs. For an index set $\mathbf{I}$, if the Bernoulli RFSs satisfy $\mathbf{X}^{i}\cap\mathbf{X}^{j}=\emptyset$ for all $i,j\in\mathbf{I}$, then the density for MB RFS $\mathbf{X}=\cup_{i\in\mathbf{I}}\mathbf{X}^{i}$ can be expressed as
@@
where $\uplus_{i\in\mathbf{I}}\mathbf{X}^{i}=\mathbf{X}$ represents all mutually disjoint (possibly empty) subsets $\mathbf{X}^{i}$, $i\in\mathbf{I}$ whose union is $\mathbf{X}$. Finally, the multi-Bernoulli mixture (MBM) density is a weighted sum of MB densities, i.e.,
@@
$f_{k|k}(\mathbf{X}_{k}|\mathbf{Z}^{k})=\sum_{\mathbf{X}^{u}_{k}\uplus\mathbf{X}^{d}_{k}=\mathbf{X}_{k}}f^{u}_{k|k}(\mathbf{X}_{k}^{u}|\mathbf{Z}^{k})f^{d}_{k|k}(\mathbf{X}_{k}^{d}|\mathbf{Z}^{k}).$ (5)
Here, $f^{u}_{k|k}(\mathbf{X}_{k}^{u}|\mathbf{Z}^{k})$ is a PPP density with intensity function $D_{k|k}^{u}(\mathbf{x})$. The MBM density $f^{d}_{k|k}(\mathbf{X}_{k}^{d}|\mathbf{Z}^{k})$ has $|\mathbf{J}_{k}|$ weighted MB densities, each of which denotes a unique history of data associations for all detected targets called global hypothesis. The $j$-th MB density $f_{k|k}^{j}(\mathbf{X}_{k}^{d}|\mathbf{Z}^{k})$ with weight $C_{k|k}^{j}$ contains $|\mathbf{I}_{k}^{j}|$ Bernoulli components, and the $i$-th Bernoulli component is determined by the existence probability $r^{j,i}_{k|k}$ and the spatial density $f_{k|k}^{j,i}(\mathbf{x})$. To sum up, the posterior PMBM density can be fully defined by a set of parameters
$D_{k|k}^{u},\{(C_{k|k}^{j},\{(r_{k|k}^{j,i},f_{k|k}^{j,i})\}_{i\in\mathbf{I}_{k}^{j}})\}_{j\in\mathbf{J}_{k}},$ (6)
@@ II-B Target State Transition Model
Given the PMBM posterior density in (5), the set of targets is partitioned into two disjoint subsets as $\mathbf{X}_{k}=\mathbf{X}_{k}^{u}\cup\mathbf{X}_{k}^{d}$, where $\mathbf{X}_{k}^{u}$ is the set of undetected targets and $\mathbf{X}_{k}^{d}$ is the set of detected targets. We assume that each target evolves with the same single-target state transition model independently over time, and new targets appear independently of the existing ones. A PPP RFS with intensity $D^{b}(\mathbf{x})$ is used to model the target birth. At time $k$, the state of a single extended target $\mathbf{x}_{k}$ is defined by a Poisson random matrix model, i.e. a combination of the kinematic vector $\mathbf{m}_{k}$, the extent matrix $E_{k}$, and the measurement rate $\gamma_{k}$. The probability of the target surviving from time step $k$ to $k+1$ is denoted by $p_{\mathrm{S}}(\mathbf{x}_{k})$.
The kinematic state (e.g., position and velocity) of the target is described by the vector $\mathbf{m}_{k}$. The target dynamic model is $\mathbf{m}_{k+1}=F_{k}\mathbf{m}_{k}+G_{k}\mathbf{w}_{k}$, where $F_{k}$ is the state transition matrix, $\mathbf{w}_{k}$ is the zero-mean Gaussian process noise with covariance matrix $Q_{k}$, and $G_{k}$ is a coefficient matrix. For the constant-turn (CT) motion model (see [17, Section V.B] for details), the kinematic state $\mathbf{m}_{k}=[x_{k},\dot{x}_{k},y_{k},\dot{y}_{k},\omega_{k}]^{\mathrm{T}}$ includes position, velocity, and turn-rate of the target. Thus, the kinematic state transition model is now defined by
@@
$G_{k}=\mathrm{diag}(G^{p},G^{p},1),\ \ G^{p}=\left[T^{2}/2,T,1\right]^{\mathrm{T}}$ (7c)
where $\mathrm{diag}(\cdot)$ forms a diagonal matrix, $T$ is the time interval, $F_{\mathrm{CT}}(\omega_{k},T)$ is the CT state transition matrix as defined in [17, (62)], $\mathrm{cov}(\cdot)$ is the covariance operation.
@@ II-C Target Measurement Model
At time $k$, the measurement set is modeled as the union of two independent subsets $\mathbf{Z}_{k}=\mathbf{Z}_{k}^{c}\cup\mathbf{Z}_{k}^{t}$, where $\mathbf{Z}_{k}^{c}$ is the set of clutters and $\mathbf{Z}_{k}^{t}$ is the set of target-originated measurements. We assume that most of the clutter measurements generated from ground surface and background objects have been removed by point cloud segmentation methods before tracking (e.g., [19, 20]). Therefore, the remaining clutters can be modeled by a PPP RFS with intensity $D^{c}(\mathbf{z})=\mu^{c}f^{c}(\mathbf{z})$.
@@
Assume that a LiDAR measurement $\mathbf{z}_{k}$ is generated from a reflective center $\mathbf{z}_{k}^{\ast}$ randomly distributed over these regions. The measurement model is defined as $\mathbf{z}_{k}=\mathbf{z}_{k}^{\ast}+\mathbf{v}_{k}$, where $\mathbf{v}_{k}=[v_{k}^{x},v_{k}^{y}]^{\mathrm{T}}$ is zero-mean Gaussian measurement noise with covariance matrix $R_{k}$. Note that $\mathbf{v}_{k}$ is an approximation of the zero-mean Gaussian measurement noise in the polar coordinate system, i.e., $\bar{\mathbf{v}}_{k}=[v_{k}^{\theta},v_{k}^{r}]^{\mathrm{T}}$ with covariance $\bar{R}_{k}=\mathrm{diag}([\sigma_{\theta}^{2},\sigma_{r}^{2}])$. Unscented transformation (UT) [21] can be used to calculate the approximate covariance $R_{k}$.
@@
where $\mathrm{mod}$ is the modulo operation, $s$ is a random scaling factor uniformly distributed over $[0,1)$. Thus, the conditional likelihood of a measurement $\mathbf{z}_{k}$ on edge region $r_{1}$ given the target state $\mathbf{x}_{k}$ can be written as
@@
Here, we assume the measurement noise over region $r_{1}$ has a static covariance matrix $\hat{R}_{k}^{1}$ approximated by UT on the center point of $r_{1}$. The integral in (II-C) is known as the stick model in literature, which is analytical resolvable [22, 23]. The measurement likelihood on the other edge regions can be calculated with formulations similar to (II-C).
@@
where $Q(\cdot)$ denotes the Q-function, $\sigma_{1}^{2}=(\mathbf{e}_{k}^{1})^{\mathrm{T}}\hat{R}_{k}^{5}\mathbf{e}_{k}^{1}$ and $\sigma_{2}^{2}=(\mathbf{e}_{k}^{2})^{\mathrm{T}}\hat{R}_{k}^{5}\mathbf{e}_{k}^{2}$ are the variances of projected measurement noise, and $d_{\mathbf{z}}^{r_{1}}$-$d_{\mathbf{z}}^{r_{4}}$ are the distances from $\mathbf{z}_{k}$ to the lines determined by edge regions $r_{1}$-$r_{4}$.
For calculating the measurement likelihoods, the continuous integrals (II-C) and (16) in our system model are more robust and accurate compared with the DRA methods, which use a finite number of random scatter centers in the calculation (see, e.g., [22][23] for related discussions).

### v3

@@ Abstract
Multiple extended target tracking (ETT) has gained increasing attention due to the development of high-precision LiDAR and radar sensors in automotive applications. For LiDAR point cloud-based vehicle tracking, this paper presents a probabilistic measurement-region association (PMRA) ETT model, which can describe the complex measurement distribution by partitioning the target extent into different regions. The PMRA model overcomes the drawbacks of previous data-region association (DRA) models by eliminating the approximation error of constrained estimation and using continuous integrals to more reliably calculate the association probabilities. Furthermore, the PMRA model is integrated with the Poisson multi-Bernoulli mixture (PMBM) filter for tracking multiple vehicles. Simulation results illustrate the superior estimation accuracy of the proposed PMRA-PMBM filter in terms of both the positions and extents of vehicles compared with PMBM filters using the gamma Gaussian inverse Wishart and DRA implementations.
@@ I Introduction
LiDAR and radar point clouds can provide abundant and accurate spatial information of the surrounding environment, which is vital for perception tasks such as target detection and tracking in autonomous driving and intelligent transportation systems [1, 2, 3, 4, 5]. In the context of point cloud-based multiple target tracking (MTT), extended target tracking (ETT) methods have attracted increasing attention [6, 7, 8]. The ETT differs from traditional MTT approaches by assuming that the sensor can gather multiple measurements of a target in each scan, thus allowing for the simultaneous estimation of the target’s location and extent directly from the point cloud [9, 10].
@@
The objective of this study is to overcome the deficiencies of DRA methods and extend the region-partitioning idea to LiDAR point cloud-based multiple vehicle tracking. To this end, we design the probabilistic measurement-region association (PMRA) model, which can directly obtain rectangular extents using the Wishart distribution. The PMRA model also utilizes continuous integrals and the visible angle of regions to improve the accuracy and stability for calculating the association probabilities. For tracking multiple vehicles, the PMRA model is integrated with the Poisson multi-Bernoulli mixture (PMBM) filter, a state-of-the-art MTT framework [15, 16]. Simulation results show that the PMRA-PMBM filter with particle implementation achieves higher estimation accuracy for both the positions and extents of vehicles compared to PMBM filters using the GGIW and DRA models.
@@ II System Modeling
A system model based on the random finite set (RFS) and the PMBM conjugate prior is applied in this study. At time step $k$, the multi-target state is represented by an RFS $\mathbf{X}_{k}=\{\mathbf{x}_{k}^{i}\}_{i\in\mathbf{I}_{k}}$ where $\mathbf{I}_{k}$ is the target index set with cardinality $|\mathbf{I}_{k}|=N_{k}$; both the single target state $\mathbf{x}_{k}^{i}$ and the number of targets $N_{k}$ are random. Similarly, the set of measurements is expressed as $\mathbf{Z}_{k}=\{\mathbf{z}_{k}^{m}\}_{m\in\mathbf{M}_{k}}$ with $|\mathbf{M}_{k}|=M_{k}$, and the collection of all measurement sets from time $1$ to time $k$ is denoted by $\mathbf{Z}^{k}$. For LiDAR-based ETT, the 3D point cloud is usually projected to the 2D bird’s-eye-view (BEV) plane before tracking [6, 7]. In this work, we follow the same approach and define the system model on the global Cartesian coordinate system, as illustrated in Fig. 1.
Figure 1: Illustration of the system model. The positions of the LiDAR and the target are denoted by $(x_{\mathrm{s}},y_{\mathrm{s}})$ and $(x_{\mathrm{t}},y_{\mathrm{t}})$. The target extent is divided into five regions $r_{1}$-$r_{5}$. The vertices $\mathbf{p}^{1}$-$\mathbf{p}^{4}$ are determined by the eigenvectors $\{\mathbf{e}^{1},\mathbf{e}^{2}\}$ and eigenvalues $\{e^{1},e^{2}\}$ of the extent matrix $E$.
@@ II-A RFSs and PMBM Posterior Density
The RFSs used in our system modeling are briefly introduced in this section. A Poisson point process (PPP) is an RFS with Poisson distributed cardinality and i.i.d. elements. Thus, the density of a PPP RFS is defined by
@@
where $\mu$ is the Poisson rate, $f(\mathbf{x})$ is the spatial density, $D(\mathbf{x})=\mu f(\mathbf{x})$ is the intensity function. A Bernoulli RFS contains a single element with probability $r$ or, is an empty set with probability $1-r$. The density of a Bernoulli RFS is
@@
A multi-Bernoulli (MB) RFS is the union of a fixed number of independent Bernoulli RFSs. For an index set $\mathbf{I}$, if the Bernoulli RFSs satisfy $\mathbf{X}^{i}\cap\mathbf{X}^{j}=\emptyset$ for all $i,j\in\mathbf{I}$, then the density for the MB RFS $\mathbf{X}=\cup_{i\in\mathbf{I}}\mathbf{X}^{i}$ can be expressed as
@@
where $\uplus_{i\in\mathbf{I}}\mathbf{X}^{i}=\mathbf{X}$ represents all mutually disjoint (possibly empty) subsets $\{\mathbf{X}^{i}\}_{i\in\mathbf{I}}$ whose union is $\mathbf{X}$. Finally, the multi-Bernoulli mixture (MBM) density is a weighted sum of MB densities, i.e.,
@@
$f_{k|k}(\mathbf{X}_{k}|\mathbf{Z}^{k})=\sum_{\mathbf{X}^{\mathrm{u}}_{k}\uplus\mathbf{X}^{\mathrm{d}}_{k}=\mathbf{X}_{k}}f^{\mathrm{u}}_{k|k}(\mathbf{X}_{k}^{\mathrm{u}}|\mathbf{Z}^{k})f^{\mathrm{d}}_{k|k}(\mathbf{X}_{k}^{\mathrm{d}}|\mathbf{Z}^{k}).$ (5)
Here, $f^{\mathrm{u}}_{k|k}(\mathbf{X}_{k}^{\mathrm{u}}|\mathbf{Z}^{k})$ is a PPP density with intensity function $D_{k|k}^{\mathrm{u}}(\mathbf{x})$. The MBM density $f^{\mathrm{d}}_{k|k}(\mathbf{X}_{k}^{\mathrm{d}}|\mathbf{Z}^{k})$ has $|\mathbf{J}_{k}|$ weighted MB densities, each of which denotes a unique history of data associations for all detected targets called the global hypothesis. The $j$-th MB density $f_{k|k}^{j}(\mathbf{X}_{k}^{\mathrm{d}}|\mathbf{Z}^{k})$ with weight $C_{k|k}^{j}$ contains $|\mathbf{I}_{k}^{j}|$ Bernoulli components, and the $i$-th Bernoulli component is determined by the existence probability $r^{j,i}_{k|k}$ and the spatial density $f_{k|k}^{j,i}(\mathbf{x})$. To sum up, the posterior PMBM density can be fully defined by a set of parameters
$D_{k|k}^{\mathrm{u}},\{(C_{k|k}^{j},\{(r_{k|k}^{j,i},f_{k|k}^{j,i})\}_{i\in\mathbf{I}_{k}^{j}})\}_{j\in\mathbf{J}_{k}},$ (6)
@@ II-B Target State Transition Model
Given the PMBM posterior density in (5), the set of targets is partitioned into two disjoint subsets as $\mathbf{X}_{k}=\mathbf{X}_{k}^{\mathrm{u}}\cup\mathbf{X}_{k}^{\mathrm{d}}$, where $\mathbf{X}_{k}^{\mathrm{u}}$ is the set of undetected targets and $\mathbf{X}_{k}^{\mathrm{d}}$ is the set of detected targets. We assume that each target evolves with the same single-target state transition model independently over time, and new targets appear independently of the existing ones. A PPP RFS with intensity $D^{\mathrm{b}}(\mathbf{x})$ is used to model the target birth. At time $k$, the state of a single extended target $\mathbf{x}_{k}$ is defined by a Poisson random matrix model, i.e. a combination of the kinematic vector $\mathbf{m}_{k}$, the extent matrix $E_{k}$, and the measurement rate $\gamma_{k}$. The probability of the target surviving from time step $k$ to $k+1$ is denoted by $p_{\mathrm{S}}(\mathbf{x}_{k})$.
The kinematic state (e.g., position and velocity) of the target is described by the vector $\mathbf{m}_{k}$. The target dynamic model is $\mathbf{m}_{k+1}=F_{k}\mathbf{m}_{k}+G_{k}\mathbf{w}_{k}$, where $F_{k}$ is the state transition matrix, $\mathbf{w}_{k}$ is the zero-mean Gaussian process noise with covariance matrix $Q_{k}$, and $G_{k}$ is a coefficient matrix. For the constant-turn (CT) motion model (see [17, Section V.B] for details), the kinematic state $\mathbf{m}_{k}=[x_{k},\dot{x}_{k},y_{k},\dot{y}_{k},\omega_{k}]^{\mathrm{T}}$ includes the position, velocity, and turn-rate of a target. Thus, the kinematic state transition model is now defined by
@@
$G_{k}=\mathrm{diag}(G^{p},G^{p},T),\ \ G^{p}=\left[T^{2}/2,T\right]^{\mathrm{T}}$ (7c)
where $\mathrm{diag}(\cdot)$ forms a block diagonal matrix, $T$ is the time interval, $F_{\mathrm{CT}}(\omega_{k},T)$ is the CT state transition matrix as defined in [17, (62)], $\mathrm{cov}(\cdot)$ is the covariance operation.
@@ II-C Target Measurement Model
At time $k$, the measurement set is modeled as the union of two independent subsets $\mathbf{Z}_{k}=\mathbf{Z}_{k}^{\mathrm{c}}\cup\mathbf{Z}_{k}^{\mathrm{t}}$, where $\mathbf{Z}_{k}^{\mathrm{c}}$ is the set of clutters and $\mathbf{Z}_{k}^{\mathrm{t}}$ is the set of target-originated measurements. We assume that most of the clutter measurements generated from the ground surface and background objects have been removed by point cloud segmentation methods before tracking (e.g., [19, 20]). Therefore, the remaining clutters can be modeled by a PPP RFS with intensity $D^{\mathrm{c}}(\mathbf{z})=\mu^{\mathrm{c}}f^{\mathrm{c}}(\mathbf{z})$.
@@
Assume that a LiDAR measurement $\mathbf{z}_{k}$ is generated from a reflective center $\mathbf{z}_{k}^{\ast}$ randomly distributed over these regions. The measurement model is defined as $\mathbf{z}_{k}=\mathbf{z}_{k}^{\ast}+\mathbf{v}_{k}$, where $\mathbf{v}_{k}=[v_{k}^{x},v_{k}^{y}]^{\mathrm{T}}$ is the zero-mean Gaussian measurement noise with covariance matrix $R_{k}$. Note that $\mathbf{v}_{k}$ is an approximation of the zero-mean Gaussian measurement noise in the polar coordinate system, i.e., $\bar{\mathbf{v}}_{k}=[v_{k}^{\theta},v_{k}^{r}]^{\mathrm{T}}$ with covariance $\bar{R}_{k}=\mathrm{diag}([\sigma_{\theta}^{2},\sigma_{r}^{2}])$. Unscented transformation (UT) [21] can be used to calculate the approximate covariance $R_{k}$.
@@
where $\mathrm{mod}$ is the modulo operation, $s$ is a random scaling factor uniformly distributed over $[0,1)$. Thus, the conditional likelihood of a measurement $\mathbf{z}_{k}$ on the edge region $r_{1}$ given the target state $\mathbf{x}_{k}$ can be written as
@@
Here, we assume the measurement noise over region $r_{1}$ has a static covariance matrix $\hat{R}_{k}^{1}$ approximated by UT on the center point of $r_{1}$. The integral in (15) is known as the stick model in literature, which is resolvable with approximations or special functions [22, 23]. The measurement likelihood formulations on the other edge regions are similar to (15).
@@
where $Q(\cdot)$ denotes the Q-function, $\sigma_{1}^{2}=(\mathbf{e}_{k}^{1})^{\mathrm{T}}\hat{R}_{k}^{5}\mathbf{e}_{k}^{1}$ and $\sigma_{2}^{2}=(\mathbf{e}_{k}^{2})^{\mathrm{T}}\hat{R}_{k}^{5}\mathbf{e}_{k}^{2}$ are the variances of the projected measurement noise, and $d_{\mathbf{z}}^{r_{1}}$-$d_{\mathbf{z}}^{r_{4}}$ are the distances from $\mathbf{z}_{k}$ to the lines determined by edge regions $r_{1}$-$r_{4}$.
For calculating the measurement likelihoods, the continuous integrals (15) and (16) in our system model are more robust and accurate compared with the DRA methods, which use a finite number of random scatter centers in the calculation (see, e.g., [22] and [23] for related discussions).